A new ad campaign came out that boosts demand. The inverse demand function is P = 108 - 1.2 Q AND supply is P = 0.5Q + 40.

Find the equilibrium point. What is the price there?

What is the quantity there?

What is the total surplus there?
Answer #1

Answer
The equilibrium is at Qd=Qs
equating the demand and supply equations
108-1.2Q=0.5Q+40
1.7Q=68
Q=40 units
P=0.5*40+40=$60

the price is $60, and the quantity is 40 units
A total surplus is between the demand and supply curve and left of the quantity.

the total surplus =0.5*(Y-intercept of demand curve -Y-intercept of supply curve)*Q
=0.5*(108-40)*40
=1360

the total surplus is $1360